OpenStudy (anonymous):

Verify the identity. cos (x - y) - cos (x + y) = 2 sin x sin y

OpenStudy (anonymous):

cos (x - y) - cos (x + y) = 2 sin x sin y LHS =cos (x - y) - cos (x + y) = cosx cosy+sinxsiny -(cosxcosy-sinxsiny) = cosx cosy+sinxsiny -cosxcosy+sinxsiny = sinxsiny +sinxsiny = 2sinx siny =RHS Thus given identity is verified.

OpenStudy (anonymous):

Gee, thanks!